OpenStudy (anonymous):

a stone is dropped from a height of 45 m what is the distance travelled by it during the last one second of motion?

OpenStudy (shamim):

Let time=t to touch the ground

OpenStudy (shamim):

Given h=45m

OpenStudy (shamim):

Initial velocity u=0

OpenStudy (shamim):

Required formula is h=ut+0.5gt^2 45=0.5*9.8*t^2 t=?

OpenStudy (shamim):

Response plz !!!

OpenStudy (shamim):

I wanna help u the rest!!